Allison:

What is the exact volume of the cylinder? 18π in³ 36π in³ 54π in³ 63π in³

Allison:

243703-VHS_PA_S2_05_L204_L304_LQ_Q1.gif

Allison:

@Hero

Hero:

@Allison do you remember the formula for volume of a cylinder? What is it?

Allison:

3.14* Radius * 2 * H

Warriorz13:

(3.14)r^2h

Warriorz13:

So, (3.14) * 3^2 * 6

Hero:

@Warriorz13, please allow @Allison to try to answer the questions.

Hero:

@Warriorz13 Allison is the one that needs help

Warriorz13:

Fine *raises hands* I'm out.

Warriorz13:

@Hero I was trying to help ._.

Hero:

Yes, but do you not understand that it is best if @Allison tries to answer the question on her own first?

Allison:

Cut it out

Hero:

It helps me figure out what she knows and doesn't know so that I can best help her.

Allison:

I got 2,513..??

Allison:

Wait

Allison:

I got 168

Hero:

Let's start over. The formula for Volume of a Cylinder is \[V_c = \pi r^2h\]

Allison:

113

Hero:

@Allison, What are the values for radius, r and height, h?

Allison:

R is 3, and H is 6.

Hero:

Very good so: \[V_c = \pi (3)^2 (6)\]

Allison:

113.04 v.v

Hero:

In other words: \[V_c = 54\pi\]

Hero:

In cubic inches of course

Allison:

o.0

Allison:

Kay. Thank you.

Warriorz13:

You take the number you get and divide it by pi or 3.14